How do you solve the equation z+6/3=8?
Found 2 solutions by stanbon, Godslove18:
Answer by stanbon(75887)

solve the equation z+6/3=8?
--------
(z+6)/3 = 8
--------
z+6 = 24
--------
z = 18
==========
Cheers,
Stan H.

Answer by Godslove18(3)

z+6/3=8 z=8-6/3 z=8(3)-6/3 z=24-6/3=18/3=6 therefore z equals 6